OpenStudy (anonymous):

Find r(t) given the following: r''(t)=cos(2t) i + sin(2t) j r'(0)=2i-1/2j r(0)=3/4i+j I know that I must integrate r''(t) twice, but is the r'(0) the constant value or what?

OpenStudy (anonymous):

yes. integrating once gives: r'(t) = (1/2)(sin(2t)i - cos(2t)j) + C taking r'(0) = 2i - (1/2)j = (1/2)(sin(0)i - cos(0)j) + C C = 2i. therefore r'(t) = (1/2)(sin(2t)i - cos(2t)j) + 2i doing this again and solving for C again will find r(t)

OpenStudy (anonymous):

is it not the constant value, but it is used to find the constant value

OpenStudy (anonymous):

Ok cool! Thanks!

OpenStudy (anonymous):

Although I'm still not sure why is C=2i when cos(0)=1 So wouldn't r'(0)=0i+2i-(1/2)j-(1/2)j?

OpenStudy (anonymous):

check again ^_^ C = 0i + 2i - (1/2)j + (1/2)J

OpenStudy (anonymous):

I'm still not sure I get it. If you simply add the given r'(0) to the found r'(0), then it is sin0-cos0, so it is (1/2)(0-1), so therefore it is (-1/2)+(-1/2)

OpenStudy (anonymous):

2i - (1/2)j = (1/2)(sin(0)i - cos(0)j) + C keep C on the right and move everything to the left. 2i - (1/2)j = (1/2)(0 - j) + C 2i - (1/2)j - (-1/2)j = C

OpenStudy (anonymous):

Awesome! I didn't understand exactly what that operation was until you wrote it out. Thanks again!